Before I started observing, the observing script website was down, after reloading and crossing fingers it went back, but there was no communication with TCS. I took the first observation manually and before the next target I re-started the Chiron GUI. For the first target T0121636363 the TCS parameters are not written in the header.
Air conditioning is not working properly. Javier is aware of the issue and will ask someone to check it.
Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
---|---|---|---|---|---|---|
2023-03-27 | 10.5 | 0.0 | 0.0 | 0.0 | 0.0 | 10.5 |
Links
[1] mailto:thenry@astro.gsu.edu
[2] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/42314/delete?destination=print/content/enr-ctio-15-m-2023-03-27
[3] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/42314/edit?destination=print/content/enr-ctio-15-m-2023-03-27
[4] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/add/node/20220?destination=print/content/enr-ctio-15-m-2023-03-27
Program completed.
DIMM didn’t work tonight, I used RINGSS data.
Javier Rojas secured the fiber.